SERIES CIRCUIT:

1. Two 80 Ohms are connected in series with 0.15A flowing through them?
FIND;
A. total resistance
B. voltage drop across resistors
C. current through the circuit

2. in this circuit, the voltage drop across the upper resistor is 4.5V.
What is the battery voltage?
What is the current through the circuit?
What is the resistance in the circuit?
(30 ohms in the upper side of the circuit and 50 ohms in the right side of the circuit)
1
Expert's answer
2021-05-24T15:48:26-0400

Gives

"R_1=R_2=80 \Omega"

Current "(i)=0.15A"



Total resistance

Series combination

"R_{net}=R_1+R_2"

"R_{net}=80+80=160\Omega"

Voltage drop across the resistance

"V_1=iR_1"

"V_2=iR_2"

"V_1=V_2=0.15\times80=12V"

Net current through the circuit

"I=\frac{V}{R_{net}}"

"I=\frac{24}{160}=0.15A"

Part (2)



Resistance of parrelled combination

"R_1=30\Omega"

"R_2=50\Omega"

Voltage drop across R1 resistance

V1=4.5V

Net resistance

"\frac{1}{R}=\frac{1}{R_1}+\frac{1}{R_2}"

"R=\frac{R_1R_2}{R_1+R_2}"


"R=\frac{30\times50}{30+50}=18.75\Omega"

Rnet=18.75 ohm

Current uper resistance

"V=i_1R_1"

"i_1=\frac{V}{R_1}"

"i_1=\frac{4.5}{30}=0.15A"


Current flow R2 resistance

"i_2=\frac{V}{R_2}"

"i_2=\frac{4.5}{50}=0.09A"


Battery voltage (v)= 4.5V

Parrelled combination voltage drop same each resistance


Current through the circuit

"I=i_1+i_2"

"I=0.15+0.09=0.24A"

"I=0.24A"